OpenStudy (anonymous):

integral (u^2)/[sqrt -4(u^2)+4]^3 du can I factor out a (-1/4) outside the integral?

myininaya (myininaya):

\[\int\limits_{}^{}\frac{u^2}{\sqrt{(-4u^2+4)^3}} du\] Is this correct?

OpenStudy (anonymous):

the sqrt is all raised to the third powr

OpenStudy (anonymous):

Yeah that is 3 only and not 2..

myininaya (myininaya):

\[\int\limits_{}^{}\frac{u^2}{(\sqrt{-4u^2+4})^3} du\]

myininaya (myininaya):

This looks like you would need a trig sub either way.

OpenStudy (anonymous):

yea, I'm prepping for a trig sub, but I want my form u^2 +4

OpenStudy (anonymous):

but since the (-4) lies within the sqrt, I'm not remembering my algebra if I can bring it out or not

myininaya (myininaya):

\[(\sqrt{-4u^2+4})^3=(\sqrt{4(-u^2+1)})^3=(\sqrt{4} \sqrt{-u^2+1})^3=(\sqrt{4})^3(\sqrt{-u^2+1})^3\] So this part right here is the part that you need help on?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

looks like rt 4^3 is a coefficient I can factor when you write it like that

myininaya (myininaya):

Yes. I just used some law of exponents stuff. Like I used \[((ab)^\frac{1}{n})^b=(a^\frac{1}{n}b^\frac{1}{n})^b=a^\frac{b}{n} b^\frac{b}{n}\]

OpenStudy (anonymous):

some tutors at school were just popping out the -4, but it didn't seem I could do that since it was within the root.

myininaya (myininaya):

You mean bring it outside the integral?

OpenStudy (anonymous):

yes

myininaya (myininaya):

Yes you can bring constant multiplies outside the integral

myininaya (myininaya):

\[\frac{1}{(\sqrt{4})^3} \int\limits_{}^{}\frac{u^2}{(\sqrt{-u^2+1})^3} du\]

OpenStudy (anonymous):

I mean that it's incorrect to go from the first equation we started with to just bringing (-1/4) outside the integral. Since (-4) is inside the sqrt, it cannot be brought outside teh integral as (-1/4).

OpenStudy (anonymous):

When you write it as 1/(sqrt 4)^3 it makes more sense.

myininaya (myininaya):

And of course that can be simplified